OpenStudy (anonymous):

Consider the functions f(x) = e^(2x) and g(x) = sin (πx/2). Find the period of the function f o g.

OpenStudy (anonymous):

same as the period of \[\sin(\frac{\pi x}{2})\]

OpenStudy (anonymous):

period of \(\sin(bx)\) is \(\frac{2\pi}{b}\)